### v2

# Towards Building Multilingual Language Model for Medicine

Abstract
In this paper, we aim to develop an open-source, multilingual language model for medicine, that the benefits a wider, linguistically diverse audience from different regions. In general, we present the contribution from the following aspects: first, for multilingual medical-specific adaptation, we construct a new multilingual medical corpus, that contains approximately 25.5B tokens encompassing 6 main languages, termed as MMedC, that enables auto-regressive training for existing general LLMs. second, to monitor the development of multilingual LLMs in medicine, we propose a new multilingual medical multi-choice question-answering benchmark with rationale, termed as MMedBench; third, we have assessed a number of popular, open-source large language models (LLMs) on our benchmark, along with those further auto-regressive trained on MMedC, as a result, our final model, termed as MMedLM 2, with only 7B parameters, achieves superior performance compared to all other open-source models, even rivaling GPT-4 on MMedBench. We will make the resources publicly available, including code, model weights, and datasets.
1 Introduction
In the recent literature, large language models (LLMs) have demonstrated great promise in healthcare, for example, closed-source models such as GPT-4 [20] and MedPalm-2 [23] have shown remarkable performance, and successfully passed the United States Medical Licensing Examination (USMLE). Concurrently, open-source models like Llama 2 have also facilitated the development of specialized language models for medicine, such as PMC-LLaMA, MedAlpaca, and ChatDoctors [29, 6, 31], gradually bridging the performance gap with their closed-source peers. Despite these advancements, the primary focus on English-language applications by these sophisticated medical language models has constrained their potential reach, limiting the benefits to a wider, linguistically diverse audience.
Figure 1: Overview of our contributions. Figure a demonstrates our proposed large-scale multilingual medical corpus (MMedC), containing 25.5B tokens, covering six main language, collected from four data sources. Figure b shows the composition our comprehensive multilingual medical benchmark (MMedBench), that is constructed by aggregating medical QA cases in different languages, and prompting GPT-4 to provide rationale sentences. MMedBench enables the evaluation on both multi-choice accuracy and the ability of rationale generation for different LLMs under zero-shot or fine-tuning settings. In Figure c left, the final multi-choice accuracy of various LLMs on our MMedBench are show, where our final model MMedLM 2 demonstrated the best performance among all existing open-source LLMs. In Figure c right, we further detail the gains in both multi-choice accuracy and ability of rationale generation, when comparing MMedLM to InternLM, or comparing MMedLM 2 to InternLM 2. Considering that the main difference between the MMedLM series and the InterLM series lies in the auto-regressive training on MMedC, such comparison highlights the importance of our contributed medical-specific multilingual language corpus.
In the realm of open-source multilingual Large Language Models (LLMs), exemplified by BLOOM [22] and the more recent InternLM 2 [25], a notable challenge persists despite their training on diverse multilingual corpora, that is, they exhibit unsatisfactory performance on medical queries in non-English languages, a discrepancy primarily attributed to the under-representation of medical content in these general datasets. This paper endeavors to bridge this gap by developing an open-source, multilingual language model for healthcare. As shown by Figure 1, our contribution is threefold: firstly, we gather a multilingual medical corpus designed for auto-regressive training, this aims to lay a robust foundation that accurately reflects the linguistic diversity and complexity of the medical domain; secondly, to monitor the progress, we introduce a new comprehensive multilingual medical question-answering (QA) benchmark, enabling the evaluation on multi-choice QA and rationale ability of different language models under both zero-shot and fine-tuning settings; lastly, we have tested a wide spectrum of existing language models, together with those that have undergone auto-regressive pre-training on our corpus. Through this comprehensive evaluation, we aim to provide valuable insights into the models’ capabilities and fostering a deeper understanding of the intricacies involved in multilingual medical query processing.
For auto-regressive training, we have developed a large-scale Multilingual Medical Corpus (MMedC), amassing over 25.5 billion medical-related tokens across six primary languages: English, Chinese, Japanese, French, Russian, and Spanish. This diverse dataset was compiled from four distinct sources: (i), we devised an automatic pipeline to filter medical-related content from the broad multilingual corpus, ensuring a focused and relevant dataset, (ii), we curated an extensive collection of medical textbooks in various languages, and converted them into texts with carefully designed pre-processing, e.g., Optical Character Recognition (OCR), heuristic data filtering, etc. Due to copyright issue, we will share the name list of the books and the methodologies and tools for curation, (iii), to guarantee a wide-ranging encapsulation of medical knowledge, we incorporated texts from some famous open-source medical websites, enriching our corpus with authoritative and comprehensive medical information, (iv), additionally, we have integrated a number of existing small-scale medical corpus datasets, further enhancing the breadth and depth of ours. To our knowledge, MMedC represents the first endeavor to construct a corpus specifically focused on the multilingual medical domain.
As for benchmark curation, we start by aggregating existing medical multiple-choice QA datasets across the six languages as of MMedC. We further augment them with rationale content by using GPT-4, enriching the datasets with explanations that support the correct answers. Consequently, our enriched dataset encompasses 53,566 QA pairs across six languages, uniquely offering both multi-choice QA and accompanying rationale reasoning. This extensive collection spans 21 medical fields, including but not limited to Internal Medicine, Biochemistry, Pharmacology, and Psychiatry, among others, termed as the Multilingual Medical Benchmark (MMedBench). We divide it into 45,048 training pairs and 8,518 testing pairs. The training split enables to finetune LLMs after domain-specific fine-tuning. We utilize the entire test set, comprising 8,518 QA pairs, to evaluate the accuracy of multi-choice question answering. To further examine the models’ reasoning ability, we select a subset of 1,136 QA pairs, each accompanied by manually verified rationale sentences, serving as a more specialized benchmark for reasoning evaluation.
At the evaluation phase, we conducted comprehensive benchmarking across eleven existing LLMs with multilingual support, including, GPT-3.5, GPT-4, Gemini-1.0 pro, BLOOM, InternLM, InternLM 2, MedAlpaca, ChatDoctor, PMC-LLaMA, and Mistral, alongside the LLMs further trained with MMedC. These models were evaluated across three different settings: zero-shot, parameter-efficient fine-tuning (PEFT), and full fine-tuning. Given the complexity of evaluating rationale quality, which demands an assessment of long sentence semantic integrity, in addition to leveraging mainstream automated metrics, we also incorporate human rating scores in our analysis. This dual approach not only provides a comprehensive measure on each model’s performance, but also enables us to scrutinize the correlation between automated metrics and human judgment. Through this analysis, we identify the most reliable metric for extended comparisons, thereby enriching the methodology for evaluating reasoning ability in large language models.
In our experiments, models that underwent further auto-regressive training on the MMedC consistently demonstrate enhanced performance, thereby underscoring the value and effectiveness of our compiled multilingual corpus. We will publicly release our dataset (except for the license-restricted books for which we will provide a names list), codebase, and trained models to foster future research. In addition, we recognize the significance of robust evaluation metrics, particularly for the generation of medical texts that often involve complex, long sentences. To this end, we will also release detailed human rating results for individual cases.
2 Results
Here, we start by presenting the statistics of our constructed datasets. Then, we evaluate various LLMs on MMedBench on multi-choice questions and rationale ability, as well as verifying the effectiveness of MMedC. Lastly, we conduct a series of ablation studies to investigate deeper on the impact of each data phase.
2.1 Data Statistics
We present the detailed statistics on the two proposed datasets, namely, MMedC, the most extensive Multilingual Medical Corpus to date, and MMedBench, a new multilingual medical benchmark.
2.1.1 Multilingual Medical Corpus (MMedC)
MMedC refers to a multilingual medical corpus, that contains over 25.5B tokens, acquired mainly from four sources, i.e., filtering medical-related content from the general large-scale multilingual corpus, medical textbooks, medical websites and existing small-scale corpus. The main statistic results are plotted in Figure 2.
In detail, our analysis begins with the composition of the Multilingual Medical Corpus (MMedC), which incorporates six languages that collectively cover a significant portion of the global population. This diversity ensures our model’s broad applicability across various linguistic contexts, as illustrated in Sub-figure (a). Subsequently, Sub-figure (b) presents a detailed breakdown of the token distribution across these languages. Notably, English constitutes the largest segment at 42%, while Russian represents the smallest at just 7%. However, it’s important to highlight that even the smallest share, given the corpus’s overall volume of 25.5 billion tokens, translates into a substantial amount of text—approximately 2 billion tokens. Lastly, Sub-figure (c) delineates the contribution of four distinct sources to our dataset across different languages. Predominantly, medically-related content filtered from broader datasets forms the bulk of contributions for most languages, supplemented by data from medical textbooks, medical websites, and pre-existing small-scale corpora. The variety of sources ensures richness of medical knowledge, ranging from everyday medical information to more specialized knowledge found in textbooks and encyclopedias. This detailed examination of our data sources sheds light on the nuanced composition of MMedC, offering insights into its diverse and comprehensive nature.
Figure 2: Statistic results on MMedC. Figure (a) shows our collected corpora can cover most main countries worldwide. Figure (b) shows the detail token number for different languages and Figure (c) shows how the four considered data sources contributes for different languages, i.e., filtering content, medical textbooks, medical websites and small-scale corpus.
2.1.2 Multilingual Medical Benchmark (MMedBench)
MMedBench represents a comprehensive multilingual medical Question and Answering (QA) benchmark. We start by delving into its core attributes, including the total number of training and testing cases, the distribution of answer options, and the average length of question-and-answer tokens. Figure 3a illustrates these fundamental characteristics, highlighting that MMedBench often includes questions with multiple correct options, that brings complexity for models to navigate. Additionally, the answers contain rationale sections averaging 200 tokens each. This substantial token count serves a dual purpose: it not only aids in training language models by exposing them to extended reasoning passages but also in evaluating their ability to generate and understand lengthy, complex reasoning statements.
In our detailed exploration of MMedBench, we categorize each question into one of 21 medical topic categories using GPT-4. These categories include Internal Medicine, Biochemistry, Pharmacology, Psychiatry, Microbiology, Physiology, Pathology, Immunology, Obstetrics and Gynecology, Public Health, Hematology, Surgery, Emergency Medicine, Orthopedics, Neurology, Anatomy, Medical Genetics, Radiology, Dermatology, and Endocrinology. This categorization has been rigorously verified by at least two clinicians to ensure its comprehensiveness, covering the breadth of medical disciplines.
Figure 3b showcases the diversity of our multilingual benchmark, spanning a wide array of medical questions from foundational clinical medicine to specialized areas such as pharmacology and public health, with a pronounced emphasis on areas like Internal Medicine and Biochemistry. This underlines the benchmark’s capacity to assess a model’s ability to recognize and process a broad spectrum of medical inquiries effectively.
Figure 3: Statistic results on MMedBench. Figure a shows the basic statistic number on the train and test set of MMedBench. The term "Avg. tokens" represents the mean token length per sample across various compositions in it. “Rationale” denotes the rationale sentences in answer. “Option” denotes the option descriptions in choice list and “question” denotes the question sentences. Then the term “Prop. of multi-option” denotes the proportion of the question with multiple correct options and “Prop. of single-option” denotes the proportion of those with one options in answer. The final term “Number of QA pairs” denotes how many QA pairs are in train or test split. Figure b shows the topics distribution in the test split of MMedBench, covering a wide range of medical aspects, ranging from general and specialized medicine to basic medical sciences. This allows MedQA to comprehensively measure the performance of medical models.
2.2 Evaluation on MMedBench
In this section, we present a comprehensive benchmark of the foremost LLMs using our MMedBench under zero-shot, PEFT, and full fine-tuning settings. Our evaluation focuses on two aspects of model performance: the accuracy in multiple-choice questions and the models’ ability to generate rationales. The evaluated LLMs can be categorized into four distinct classes, i.e., close-source LLMs, popular open-source LLMs, medical-specific open-source LLMs, and the those further undergone training on our MMedC. The later three can all be categorized into open-source LLMs.
Initially, our analysis focuses on the state-of-the-art, proprietary close-source LLMs developed by OpenAI and Google, specifically GPT-3.5, GPT-4, and Gemini-1.0 pro. These models are examined through their publicly available online API solely in a zero-shot setting, as they are not accessible for any further training. However, note that, as the training data for these close-source models are confidential, it is difficult to judge whether they are really “zero-shot”. Following this, our evaluation encompasses a range of open-source LLMs such as Llama 2, Mistral, and InternLM 2. We observe that the response from these open-source LLMs to instructions is relatively poor, making it difficult to draw effective comparisons (check Supplementary Material D for more zero-shot fail cases) in the zero-shot setting. We therefore compare them in fine-tuning settings (PEFT and full fine-tuning). Among these, we make a further distinction between general LLMs and those specifically tailored for the medical domain. Finally, we evaluate models that have undergone further training on our proposed corpus, named MMedLM (based on InternLM) and MMedLM 2 (based on InternLM 2). These models are uniquely augmented with domain-specific knowledge, achieved by auto-regressive training on the MMedC.
2.2.1 On Multilingual Multiple-choice Question & Answering
As illustrated in Table 1, medical-specific Large Language Models (LLMs) generally exhibit high accuracy scores in English, yet their performance significantly declines in languages other than English. Notably, the finetuned PMC-LLaMA achieved an English accuracy score of 47.53, despite outperforming its contemporaneous counterparts, falls behind the GPT models significantly. Later, with the deployment of more advanced foundational models, open-source models started to bridge the gap with the GPT series, for instance, Mistral and InternLM 2, once fine-tuned on the train set of MMedBench, recorded average accuracy scores of 60.73 and 58.59, respectively, surpassing all predecessors of comparable scale. Enhanced performance is also observed after additional auto-regressive training on our proprietary MMedC dataset. Specifically, both MMedLM and MMedLM 2, built on top of InternLM and InternLM 2 respectively, demonstrated significant improvements over their counterparts without training on MMedC, for example, 67.30 (MMedLM 2) vs. 58.59 (InternLM 2) under full fine-tuning evaluation. A similar story holds for the PEFT setting, i.e., later LLMs performs better and training on MMedC brings a significant gain. As a result, MMedLM 2 refers to the most competitive open-source model with a 7B parameter count in proximity to GPT-4’s accuracy of 74.27.
Table 1: Mutliple-choice accuracy evaluation on MMedBench. We report each model’s accuracy across various languages separately, with “Avg.” denoting the mean score over six languages under zero-shot, PEFT and full fine-tuning settings. We also list out their model sizes, release time and whether they are testing after further training on our MMedC or the training set of MMedBench in the table. The best results under each setting are bold.
[b]
Method Size Year MMedC MMedBench English Chinese Japanese French Russian Spanish Avg.
Zero-shot Evaluation
GPT-3.5 - 2022.12 56.88 52.29 34.63 32.48 66.36 66.06 51.47
GPT-4 - 2023.3 78.00 75.07 72.91 56.59 83.62 85.67 74.27
Gemini-1.0 pro - 2024.1 53.73 60.19 44.22 29.90 73.44 69.69 55.20
Parameter-efficient Fine-tuning Evaluation
BLOOMZ 7B 2023.5 trainset 38.88 48.86 17.59 18.65 53.91 44.78 37.11
InternLM 7B 2023.7 trainset 40.93 52.19 27.14 18.81 46.88 40.34 37.71
Llama2 7B 2023.7 trainset 37.00 37.13 24.12 19.13 63.67 42.89 37.32
ChatDoctor 7B 2023.3 trainset 36.68 34.06 28.14 11.58 60.55 39.86 35.15
MedAlpaca 7B 2023.4 trainset 43.28 36.81 27.14 16.40 51.95 41.72 36.22
PMC-LLaMA 7B 2023.4 trainset 33.62 31.76 20.60 10.13 57.81 37.89 31.97
Mistral 7B 2023.10 trainset 55.38 50.23 37.69 40.19 71.88 61.60 52.83
InternLM 2 7B 2024.2 trainset 52.40 68.18 39.20 28.78 63.67 55.25 51.25
MMedLM (Ours) 7B - trainset 41.16 52.22 27.14 18.49 47.66 40.34 37.83
MMedLM 2 (Ours) 7B - trainset 58.13 70.43 54.27 38.26 71.88 64.95 59.65
Full Fine-tuning Evaluation
BLOOMZ 7B 2023.5 trainset 43.28 58.06 32.66 26.37 62.89 47.34 45.10
InternLM 7B 2023.7 trainset 44.07 64.62 37.19 24.92 58.20 44.97 45.67
Llama 2 7B 2023.7 trainset 43.36 50.29 25.13 20.90 66.80 47.10 42.26
MedAlpaca 7B 2023.3 trainset 46.74 44.80 29.64 21.06 59.38 45.00 41.11
ChatDoctor 7B 2023.4 trainset 43.52 43.26 25.63 18.81 62.50 43.44 39.53
PMC-LLaMA 7B 2023.4 trainset 47.53 42.44 24.12 20.74 62.11 43.29 40.04
Mistral 7B 2023.10 trainset 61.74 71.10 44.72 48.71 74.22 63.86 60.73
InternLM 2 7B 2024.2 trainset 57.27 77.55 47.74 41.00 68.36 59.59 58.59
MMedLM (Ours) 7B - trainset 49.88 70.49 46.23 36.66 72.27 54.52 55.01
MMedLM 2 (Ours) 7B - trainset 61.74 80.01 61.81 52.09 80.47 67.65 67.30
Table 1: Mutliple-choice accuracy evaluation on MMedBench. We report each model’s accuracy across various languages separately, with “Avg.” denoting the mean score over six languages under zero-shot, PEFT and full fine-tuning settings. We also list out their model sizes, release time and whether they are testing after further training on our MMedC or the training set of MMedBench in the table. The best results under each setting are bold.
2.2.2 Machine Rating on Rationale
In addition to multiple-choice QA tasks, our study extends to examining the rationale ability of various LLMs. To facilitate this comparison, we employ several automatic metrics, namely BLEU [21] and ROUGE [16], which assess sentence similarity based on n-grams. Furthermore, we explore the use of BERT-score [34], a metric that uses a pre-trained BERT model to extract high-level semantic features and employs cosine similarity for semantic evaluation.
We offer detailed instructions that prompt the model to outline its analytical process for delivering the final answer, enabling a clear assessment of its reasoning capabilities. The performance is then meticulously evaluated using a variety of metrics. Specifically, the ROUGE-1 and BLEU-1 scores are presented in Table 2. Additionally, results for other metrics are detailed in Supplementary Material C providing a comprehensive view of the model’s performance across diverse evaluation frameworks.
Table 2: Rationale evaluation on MMedB with ROUGE-1/ BLEU-1. The value preceding the symbol ‘/’ represents BLEU-1, while the value following it represents ROUGE-1. The best results are bold under different settings.
[b]
Method English Chinese Japanese French Russian Spanish Avg.
Zero-Shot Evaluation
GPT-3.5 36.21/ 38.25 27.33/ 37.34 21.30/ 31.87 33.95/ 45.51 12.65/ 20.70 24.62/ 36.20 26.01/ 34.98
Gemini-1.0 pro 11.85/ 28.20 6.26 / 27.23 6.54 / 24.28 8.42/ 33.11 3.39/ 15.38 7.22/ 27.98 7.28/ 26.03
Parameter-efficient Fine-tuning Evaluation
BLOOMZ 41.45/ 36.81 43.09/ 45.17 28.79/ 38.09 38.89/ 37.49 22.25/ 15.28 42.36/ 39.22 36.14/ 35.34
InternLM 41.29/ 38.05 43.47/ 44.71 22.80/ 37.57 30.35/ 32.14 18.24/ 16.79 36.32/ 34.56 32.08/ 33.97
Llama 2 44.72/ 39.34 42.69/ 43.71 45.58/ 49.53 42.93/ 39.29 31.75/ 22.66 44.22/ 39.64 41.98/ 39.03
ChatDoctor 44.65/ 40.26 40.88/ 42.80 39.54/ 45.00 40.12/ 39.06 30.95/ 22.84 42.88/ 40.23 39.84/ 38.37
MedAlpaca 43.59/ 39.52 40.71/ 42.50 37.27/ 44.69 39.82/ 39.57 30.11/ 22.83 42.80/ 39.64 39.05/ 38.12
PMC-LLaMA 44.98/ 40.90 40.09/ 42.95 38.15/ 43.67 38.89/ 38.64 30.08/ 22.45 43.00/ 39.80 39.20/ 38.07
Mistral 48.13/ 42.80 45.61/ 46.31 43.82/ 48.19 44.73/ 41.07 33.62/ 24.75 47.37/ 42.83 43.88/ 40.99
InternLM 2 46.87/ 41.66 47.64/ 49.28 42.22/ 46.91 41.81/ 38.46 26.78/ 21.71 44.51/ 40.13 41.64/ 39.69
MMedLM 41.63/ 38.83 44.30/ 46.38 38.61/ 46.90 37.54/ 37.78 19.99/ 21.28 40.79/ 38.77 37.14/ 38.32
MMedLM 2 47.07/ 41.51 47.15/ 48.36 47.90/ 52.24 43.22/ 41.36 27.81/ 25.70 46.17/ 42.64 43.22/ 41.97
Full Fine-tuning Evaluation
BLOOMZ 45.94/ 40.51 48.37/ 48.26 44.71/ 48.61 44.47/ 41.05 29.95/ 21.50 45.91/ 40.77 43.22/ 40.12
InternLM 46.53/ 41.86 48.24/ 48.64 44.89/ 49.83 41.80/ 37.95 27.87/ 21.20 43.42/ 38.59 42.12/ 39.68
Llama 2 46.87/ 41.39 46.62/ 46.57 48.53/ 51.21 44.43/ 40.38 33.05/ 23.24 45.96/ 40.37 44.24/ 40.53
MedAlpaca 47.33/ 42.31 45.72/ 46.49 45.35/ 49.12 43.78/ 40.41 32.80/ 23.15 45.99/ 40.57 43.49/ 40.34
ChatDoctor 47.22/ 41.97 44.66/ 45.81 38.87/ 47.95 44.64/ 40.25 32.19/ 23.37 45.68/ 40.71 42.21/ 40.01
PMC-LLaMA 47.33/ 42.87 45.87/ 46.18 44.52/ 48.44 43.80/ 40.23 31.14/ 22.28 46.30/ 40.68 43.16/ 40.12
Mistral 47.16/ 41.82 48.34/ 47.91 48.80/ 50.60 45.83/ 40.88 34.52/ 24.68 47.55/ 41.41 45.37/ 41.22
InternLM2 49.48/ 44.12 51.38/ 51.58 50.64/ 53.46 46.73/ 42.00 32.93/ 24.05 47.94/ 41.96 46.52/ 42.86
MMedLM 47.37/ 41.98 48.68/ 49.28 48.95/ 52.34 45.39/ 41.41 33.24/ 24.67 46.68/ 41.35 45.05/ 41.84
MMedLM 2 50.02/ 44.77 51.39/ 51.78 54.79/ 57.10 49.04/ 45.30 37.49/ 28.18 50.14/ 44.59 48.81/ 45.29
Table 2: Rationale evaluation on MMedB with ROUGE-1/ BLEU-1. The value preceding the symbol ‘/’ represents BLEU-1, while the value following it represents ROUGE-1. The best results are bold under different settings.
2.2.3 Human Rating on Rationale.
Given the limitations of automatic metrics in evaluating free-text generation, we further employ relative human ratings to rank performance and identify the most reliable automatic metric for future in-depth evaluations.
Specifically, from the test set of MMedBench, we randomly selected 50 test cases per language, alongside outcomes generated by six notable models: our MMedLM, InternLM, BLOOM, LLaMA 2, PMC-LLaMA, and GPT-3.5. The sequence of samples and corresponding model outputs was randomized to prevent bias. The evaluation panel, comprising five post-graduate students from the medical school of Shanghai Jiao Tong University and Peking Union Medical College, was instructed to rank the outputs based on accuracy, reasoning ability, and internal knowledge. To facilitate accurate assessment, manually verified references were also provided. Rankings were quantitatively assigned, with the highest rank awarded a score of 6 and the lowest a score of 1, thereby quantifying the quality of each model’s output. In parallel, we leveraged GPT-4 as an additional evaluator, assigning it the role of a judge to rank the outputs. Further details on GPT-4’s evaluation method are available in Supplementary Material A.
Figure 4a illustrates the comparative analysis of model performances through relative ratings. Notably, MMedLM achieved the highest scores in both human (4.54) and GPT-4 (4.39) evaluations, aligning with its superior performance as indicated by the automatic machine metrics. Interestingly, GPT-3.5 received a lower human rating of 3.08, suggesting that the evaluators’ preferences might be influenced by the brevity of the responses. Comprehensive rating results for each language and model are detailed in Supplementary Material C. Given that the rating metric is a relative indicator, adding new models requires a complete re-evaluation process, which is time-consuming due to extensive human labor. Consequently, the latest LLMs such as Mistral, InternLM 2, and our MMedLM 2 have not participated in rating comparisons.
In addition to comparing different LLMs, our study delves into the correlation between various automatic evaluation metrics and human preferences. This correlation analysis enables us to identify the most effective automatic metric for benchmarking purposes, thereby potentially eliminating the need for resource-intensive human evaluations in future research. We employ the Kendall rank correlation coefficient to measure the agreement between the rankings of each model’s generated rationales by automatic metrics and human evaluations. The findings, illustrated in Figure 4b, indicate that GPT-4’s evaluation results have the highest correlation with human judgments, with a $\tau$ value of 0.555. However, it’s important to note that GPT-4’s ratings, while highly correlated, are relative and not easy to scale for evaluating newly introduced models. Among the absolute automatic metrics, ROUGE-1 emerged as the most reliable indicator, demonstrating a $\tau$ value of 0.512. For BLEU-based metrics, BLEU-1 proved to be the most predictive. Consequently, we advocate for the use of ROUGE-1 and BLEU-1 as benchmarks for assessing the rationale capabilities of newly introduced LLMs on MMedBench in subsequent studies.
Figure 4: Comparative analysis on model ratings. In Figure a, BLEU score rating denotes the rating score calculated based on ranking by BLEU score. Human rating refers to rankings provided by humans, while GPT-4 rating refers to rankings generated by GPT-4. Figure b presents the correlation between human rating results and different automatic metrics. $\tau$ is the Kendall rank correlation coefficient while $k$ is the slope of fitted line.
2.3 Ablation Studies on Data Composition
In this section, we present an analysis on the effects of data construction process, as depicted in Table 3. Our ablation studies are carried out both on MMedLM and MMedLM 2 under the full fine-tuning setting, leveraging the InternLM and InternLM 2 framework respectively. Overall, the results observed on the two models are mostly consistent, in the following, we will thus focus our discussion on the results on MMedLM 2.
Here, we distinguish HQ-Data (High-Quality Data) and US-Data (Unspecified Source Data). HQ-Data includes content sourced from books and websites, which has undergone thorough human verification, whereas US-Data is derived from filtering medical-related content from a general corpus. The outcomes, detailed in Table 3, reveal that equipping the model with comprehensive rationales results in an average multiple-choice accuracy increase of 2.42 points, elevating from 56.17 to 58.59. However, further auto-regressive training exclusively on the English segment of MMedC does not yield an overall accuracy improvement. This is primarily because the model overfits the English domain, leading to superior performance in English but shortcomings in other languages (check Supplementary Material C for more details). If expanding the auto-regressive training into the whole multilingual medical corpus, the problem will be relieved, significantly improving the final results. This includes not only boosting the choice accuracy to 51.33 but also enhancing reasoning capabilities by 2.34 and 1.54 points respectively on BLEU-1 and ROUGE-1 scores. Moreover, the inclusion of an automatically gathered US-Data facilitates an additional accuracy boost from 51.33 to 55.01, representing a significant increase of 3.68 points. Gain can also be observed in rationale ability, i.e., 0.59 in BLEU-1 and 0.47 in ROUGE-1.
Table 3: Ablation study on MMedB. We reported all ACC, BLEU and Rouge score to give an overall knowledge of the effect of each step.
[b]
Method English Multilingual HQ-Data US-Data QA Answer Rationale ACC BLEU-1 Rouge-1
MMedLM
Baseline(InternLM) 43.33 / /
Baseline(InternLM) 45.66 42.12 39.68
[…]
MMedLM 55.01 45.05 41.84
MMedLM 2
Baseline(InternLM 2) 56.17 / /
Baseline(InternLM 2) 58.59 46.52 42.86
MMedLM $\mathrm{2}_{en}$ 58.28 46.46 42.99
MMedLM $\mathrm{2}_{ocr}$ 64.43 47.95 44.57
MMedLM 2 67.30 48.81 45.29
Table 3: Ablation study on MMedB. We reported all ACC, BLEU and Rouge score to give an overall knowledge of the effect of each step.
3 Discussion
3.1 Auto-regressive Training on MMedC is Effective.
As revealed in Table 1, both MMedLM and MMedLM 2 demonstrated significant improvement over their original baseline models, namely, InternLM and InternLM 2, underscoring the effectiveness of MMedC for providing targeted domain-specific enhancements across a wide range of existing open-source language models. In addition, the observed performance boosts serve as an indication that the pre-training corpora of existing LLMs exhibit limitations when faced with multilingual medical contexts. Our findings reinforce the necessity of specialized corpora like MMedC to bridge these gaps.
3.2 Incorporating More Data is Generally Effective.
While exploring how varying data sources influence the outcomes of language model performance, our findings, presented in Table 3, reveal that, the inclusion of high-quality multilingual data (HQ-Data) can lead to significant performance improvements. Additionally, we observe that incorporating data filtered from general contexts, despite its relatively lower quality compared to more explicitly medical-related sources, is also effective. This improvement underscores the value of integrating diverse data types within MMedC, to enrich the model’s understanding of medical information across different languages.
3.3 Incorporating Rationale For Fine-tuning is Effective.
While fine-tuning on MMedBench (trainset), we observed that, the integration of rationale data with multiple-choice prediction, can enhance performance on specific tasks. As shown by Table 3, combining correct answers with their rationales during the supervised fine-tuning phase not only enables LLMs to output rationale sentence, but also results in a noteworthy multiple choice accuracy improvement of 2.33% for InternLM and 2.42% for InternLM 2 on the MMedBench (testset). This indicates that the two tasks are strongly correlated and reinforces the significance training on multi-choice prediction and rationale tasks jointly.
3.4 Strong Foundational LLMs Improve the Final Results.
On MMedBench, we also notice that stronger LLM backbones (commonly released later) generally improve the final results on multilingual medical QA. With the release of more advanced LLMs, their pre-training corpus have been expanded significantly, gradually encompassing more languages. Even though non-English languages constituted a small fraction of the total, the sheer volume of the overall corpus allowed the models to encounter a vast array of multilingual texts during training, significantly enhancing their multilingual capabilities, as seen with the comparison between LLaMA 2 and Mistral, where the later Mistral performs much better. Such enhancement in general multilingual language abilities can also improve the performance after adaptation in medical domain (MMedLM vs. MMedLM 2). This observation encourages us to focus more on building up an open-source dataset for medicine, allowing future works to better leverage the rapid improvement on general LLMs.
3.5 The Impact of Multilingual Medical LLMs.
In this work, we initialize the development of medical multilingual LLMs, which has great research and clinical impacts. In research domain, it can promote the following critical directions:
  • General Medical Artificial Intelligence (GMAI). GMAI [17] commits to developing a multimodal AI model that can be directly applied to a wide range of healthcare scenarios, where LLMs are often used as an human-machine interface [18, 30, 27]. Replacing the English-central LLM with a multilingual one enables to make good use of world-wide data source, thus expanding the available multimodal training data, and improving the representation quality for other modalities.
  • Retrieval Augmented Generation. Hallucination is considered as a major problem with currently existing LLMs, especially for medical domain. One potential solution is to develop retrieval-augmented architectures [32, 33, 15]. The key motivation is that by retrieving and inputting facts from extra knowledge bases into LLMs, the generation output can avoid most fatal fact error. However, until now, most efforts are made in English, greatly limiting retrieval augmentation methods to leverage medical knowledge in other languages. Developing multilingual LLMs can benefit the retrieval process, greatly enriching the potential available knowledge base.
On clinical practice, open-source multilingual medical LLMs can also meet the following demands:
  • Language Barriers. In many healthcare systems, language barriers between patients and healthcare providers can hinder effective communication, leading to misunderstandings, misdiagnoses, and inadequate care, causing high-quality medical resources inaccessible for most people. Multilingual medical LLMs can facilitate real-time translation and interpretation, ensuring that patients can effectively communicate their symptoms and understand their diagnoses and treatment options.
  • Cultural and Law Sensitivity. Multilingual medical LLMs can also be trained to recognize and address cultural or law nuances and sensitivities for different countries in healthcare interactions. Understanding cultural backgrounds and law differences can significantly enhance the trust to medical LLMs, leading to better health outcomes.
  • Education. These models can also be customized for education, especially in regions where there is a shortage of medical educators or resources. By providing educational materials and simulations in multiple languages, medical multilingual LLMs can help standardize medical training and ensure consistent quality of care worldwide.
3.6 Limitations and Future Work.
While our work primarily focuses on constructing a multilingual medical corpus and enhancing the capabilities of LLMs in the medical domain across various languages, we encountered certain limitations. First, there are only six languages in our dataset. In the future, we anticipate expanding to include more languages, such as German and Arabic. Second, considering computational costs, our final model is in 7B scale, we might switch the training progress to a larger architecture that may achieve better results according to the scaling law [11]. Lastly, our final model shares the hallucination flaw with other LLMs. In the future, we aim to investigate more retrieval augmentation methods [32].
4 Methodology
In this part, we will introduce our methodology in detail. Specifically, in Sec. 4.1, we introduce the construction pipeline for MMedC. In Sect. 4.2, we describe the auto-regressive training procedure. In Sect. 4.3, we discuss the new multilingual medical benchmark, MMedBench, including its curation procedure, evaluation settings and metrics.
4.1 Large-scale Multilingual Medical Corpus
We herein develop a new large-scale multilingual medical corpus, MMedC, to help enrich LLMs with domain-specific medical knowledge across different languages. In detail, we explore four primary sources, e.g., filtering medical-related content from general language corpus, medical textbooks, open-source medical websites and existing small-scale multilingual medical corpus. As a result, MMedC contains over 25B tokens, covering 6 main languages, .e.g, English, Chinese, Japanese, French, Russian, and Spanish. Next, we will introduce the data collection process from the four sources respectively.
4.1.1 Filtering Medical-related Content
In the broader landscape of natural language processing, the general NLP community has amassed an extensive array of corpora, such as CommonCrawl, which captures billions of web pages monthly and has been operational for years. Despite medical-related content only constitutes a small fraction of this colossal dataset, its sheer volume presents a valuable opportunity for creating a large-scale, medical-specific corpus with the application of sophisticated auto-filtering techniques.
Our methodology begins with the CulturaX dataset [19], a meticulously curated multilingual version of CommonCrawl, with 6.3 trillion tokens. We first introduce a rule-based filtering pipeline to sift through this dataset for medical content. This process involves the careful selection of 200 medically relevant terms per language, encompassing fields such as medicine, pharmacy, and medical biology. Given the space limitation in the paper, we will list all 1200 terms in our GitHub repositories. For sentences that utilize spaces for word separation, our approach includes word segmentation followed by keyword matching. Conversely, for sentences without clear word demarcation, we employ direct keyword matching. Utilizing the matching results, we establish two principal metrics:
Algorithm 1 Determining Medical-Related Text Samples
Input: Text $T$, Set of keywords $K$, Language type $Lang$
Output: True or False
Define $T_{C}$ as threshold for $MKC$ and $T_{D}$ for $DENS$
if $Lang$ = "Space Delimited" thennormal-$\triangleright$ Split text into words first for space delimited languages
     Segment $T$ into words based on spaces
end if
Initialize $K_{U}=\emptyset$
Initialize total keyword length $totalLen\leftarrow 0$
for each word $t$ in $T$ do
     if $t\in K$ then
         Increment $totalLen$ by $len(t)$
         if $t\notin K_{U}$ then
              Add $t$ to $K_{U}$
         end if
     end if
end for
Calculate $MKC$ and $DENS$
if $MKC>T_{C}$ and $DENS>T_{D}$ then
     return True $\triangleright$ Text is considered medical-related
else
     return False $\triangleright$ Text is not considered medical-related
end if
Medical Keyword Count. Medical keyword count quantifies the number of unique medical keywords in the texts. Let $K$ be the set of a priori keywords representing medical terms of interest, and let $T$ denote the entire text corpus under analysis. The set of unique keywords appearing in the text can be formulated as $K_{U}=\{k\mid k\in T\wedge k\in K\}$. The Medical Keyword Count (MKC) is then defined as $\text{MKC}=|K_{U}|$, where ||$\cdot$|| ⋅ | denotes the cardinality of the set.
Keyword Density. Keyword density measures the proportion of text occupied by medical keywords relative to the total text length. This metric is instrumental in identifying texts that, despite their length, only incidentally include medical terms. Let $len(T)$ denote the total number of characters in the text $T$, and $occ(t,T)$ denote the occurrence times of word $t$ in $T$. The keyword density, denoted as $D$, can be formulated as:
$D=\frac{\sum_{k\in K}len(k)\cdot occ(k,T)}{len(T)}$
With the two metrics, we simply set a threshold bar to filter each sentence. To control the filtering quality, we randomly sampled $100$ sentences per language, and on average, 98 sentences are manually checked as medical-related. The final threshold and filtering ratios are detailed in Supplementary Material B.
4.1.2 Medical Textbooks
In addition to filtering the general language corpus, we also collect dozens of medical textbooks, that represent a rich repository of extensive medical knowledge, underscored by a rigorous publication process that ensures content quality. We have curated a collection exceeding 20,000 books, in line with the methodology outlined in PMC-LLaMA [29]. To extract texts from the books, we adopted Optical Character Recognition (OCR) models, specifically, we used the PaddleOCR tool for its proficiency in handling multiple languages. The OCR process generates a list detailing the coordinates and content of each text box, which is then organized in a left-to-right and top-to-bottom order. Furthermore, to ensure a focus on medical content, we excluded non-essential pages, such as covers, tables of contents, and epilogues, identifying them by their page numbers for removal. Quantitatively, we finally collect 4B tokens for English, 1.1B tokens for Chinese, 0.4B tokens for Russian and 0.3B tokens for French.
4.1.3 Medical Websites
Considering that filtering-based data are based on CommonCrawl, which is randomly scratched and untraceable, to avoid missing some important medical knowledge websites, we further crawled a number medical-related websites as compensatory. We focus on three types of websites. Firstly, we target medical encyclopedias, which offer detailed information on diseases and drugs. While this data is of exceptional quality, it is often limited in quantity and subject to stringent access controls. Secondly, we source content from medical consultation platforms and popular science articles about medicine. These sources, though less technical, provide a wealth of knowledge on medical common sense. Lastly, we expand our data collection to include medical news websites, which allow us to gather a larger volume of unrestricted data and incorporate timely information into our model. This strategy enhanced the model’s ability to understand and respond to current medical events and trends. Collecting data from these varied websites, we compile a comprehensive and diverse medical corpus, encompassing in-depth professional medical knowledge as well as a broad spectrum of general medical information and up-to-date industry insights. The detailed websites are listed and as a result, we get 0.1B tokens for Japanese, 0.05B tokens for Spanish, and 0.1M tokens for French.
4.1.4 Existing Small-scale Multilingual Medical Corpus
Apart from the above newly collected data, we also leverage many existing open-source corpus. Specifically, we utilized the following three datasets: Wikipedia [5], Baidu Baike [3], and UFAL Medical Corpus [8]. For Wikipedia and Baidu Baike, we employ the same filtering methodology outlined in Section 4.1.1 to extract the medical domain corpus, while for UFAL, a medical corpus designed for translation tasks, we use it as it is.
4.2 Auto-regressive Training on MMedC
Once constructed MMedC, we further pre-train the existing LLMs on it in an auto-regressive manner. We adopt loss for next token prediction as used in GPT [20]. Specifically, we treat medical text as a sequence of tokens, denoted as $X=\{x_{1},x_{2},\dots,x_{N}\}$, where each $x_{i}$ is a text token and $N$ represents the total length of the sequence. For a token $x_{i}$ to be predicted, the optimization objective is:
$L(\phi)=-\sum\log(\Phi(x_{i}|x_{<i}))$ (1)
4.3 Comprehensive Multilingual Medical Benchmark
In addition to the multilingual datasets for training, we also collect a comprehensive Multilingual Medical Benchmark, spanning 6 principal languages, namely MMedBench, to conduct a thorough evaluation of a model’s performance in the medical domain across diverse languages Specifically, we started by collecting existing medical question-answering (QA) benchmarks for each language, and expand these multi-choice QA with corresponding explanations using GPT-4, followed by strict human verification to ensure the correctness of contents.
4.3.1 Multilingual Medical QA Dataset
Figure 5: The pipeline of MMedB construction. Firstly multi-choice QA pairs from various languages are collected from 5 QA datasets. Then corresponding rationale is generated with the help of GPT4. The rationale of testset is further checked by humans to ensure its quality.
Evaluating the performance of Large Language Models (LLMs) has conventionally relied on the utilization of multiple-choice questions. This evaluation framework presents the question and its corresponding options to the model, which is then expected to identify the correct answer’s index. Accuracy serves as the primary quantitative metric in this method, providing a direct and objective measure of the performance. Despite its efficacy, the prevalent medical multi-choice QA benchmarks are exclusively monolingual, thus falling short of adequately assessing LLMs’ capabilities across diverse languages. To address this deficiency and foster a more inclusive evaluation landscape, our approach involves the aggregation of various medical multi-choice QA datasets from multiple languages. This initiative aims to compile a comprehensive benchmark that reflects the multilingual realities of the medical field. The following benchmarks are considered:
  • MedQA [10] is a collection of medical multiple-choice questions (each with four answer options) based on the USMLE examination. It encompasses data in three languages: English, Simplified Chinese, and Traditional Chinese. For our evaluation, we exclusively utilize the English and Simplified Chinese sections. The data is partitioned by official guidelines.
  • lgakuQA [12] is a Japanese medical multi-choice question dataset, which comes from Japanese medical licensing examinations from the past five years (2018-2022). Since there is no official data division, we randomly divide the data and get 1,590 training samples, 199 validation samples, and 199 test samples.
  • FrenchMedMCQA [14] is a French medical multi-choice question dataset, which comes from real exams of the French medical specialization diploma in pharmacy. The data is divided according to the official release.
  • RuMedDaNet [4] is a Russian medical judgment question dataset, which we process into a binary-choice question format. The data is divided according to the official way.
  • Head-QA [28] is a Spanish multiple-choice question dataset. Questions come from exams to access a specialized position in the Spanish healthcare system. The data is divided according to the official way.
As a result, we collect 53566 QA pairs in total, for those without an official definition of train-test sets, we split them in 8:1:1 for training, validating, and testing respectively, resulting in 45048 pairs for training and 8518 for testing.
4.3.2 Rationale Generation
While the accuracy of multi-choice question answering is a straightforward and accurate metric, it fails to evaluate the reasoning and long sentence generation abilities of LLMs, which is critical for clinical usage. Therefore, we further augment each question with a justification for selecting the correct option. At evaluation phase, we prompt the model to articulate the rationale behind its choice, thereby offering insights into the model’s reasoning capabilities, as shown by Figure 5.
In detail, given GPT-4’s demonstrated capability to outperform human experts by providing detailed explanations in Chain of Thought (CoT) experiments [13], we utilize GPT-4 to generate rationales for our dataset. To guarantee the quality of these explanations, we subsequently perform human verification. Specifically, we input the question, the options, and the correct choice into GPT-4, instructing it to generate a detailed rationale for selecting a particular option. The instructions are as follows, where “{language}” will be replaced by a certain language name, like Chinese or French, in a certain case:
You’re a {language} doctor. Analyze the reasons behind choosing this particular option in 100 words for the following question in {language}.
Following the rationale generation by GPT-4, we conducted manual review to assess their quality. Our evaluation criteria are two-fold: firstly, the explanation provided by GPT-4 had to be consistent with the established correct answer for the question; secondly, it’s required to articulate the logic underpinning the answer, rather than merely replicating it. Note that, considering the cost of human checks, this will only be performed on part of our test set. Specifically, we randomly selected 200 samples from the former test split for each language to form a new rationale split. Then we shuffled and distributed among three annotators for manual verification. Annotators were tasked with categorizing each rationale as either qualified or unqualified based on the aforementioned criteria. Remarkably, we observed that 94.7% of the rationales generated by GPT-4 adhered to our standards, underscoring the high quality of the explanations. During the final evaluation phase, the calculation of rationale similarity was exclusively applied to these human-verified passed samples.
Finally, we get 1136 human-checked samples for rationale evaluation and compensate the former split 45048 training QA pairs with auto-generated rationale sentences. Given a language model, it can use our training set to further fine-tune and then be evaluated or directly evaluated on the rationale and choice testing sets.
4.3.3 Topic classification
Here, we explore the thematic distribution of the samples. For this purpose, we employ GPT-4 to categorize the topics within the test set. The instructions provided to GPT-4 for topic classification are outlined as follows, where similarly, “{language}” will be replaced by a certain language name.: You’re a {language} doctor, choose one subject out of {medical_subjects_string}, which is most relevant to the following question. At times, GPT-4 may yield ambiguous classification outcomes, such as those not aligning with predefined medical subjects. In these instances, we prompt GPT-4 to try the classification up to four times. Should it fail to produce a categorization that adheres to our criteria, we assign the sample’s category as ‘None’. Given the rarity of such occurrences, their impact on the overall statistics is minimal.
4.3.4 Evaluation Settings
To comprehensively assess the model’s performance, we tested it in three different evaluation settings: zero-shot, parameter efficient fine-tuning (PEFT), and full fine-tuning. For the zero-shot setting, we directly test off-the-shelf LLMs with proper instruction, without any further exposure to the training part of MMedBench. In addition to zero-shot, to better evaluate the performance differences between models, we also try to fine-tune and then test the open-source models. There are two mainly used fine-tuning approaches: parameter efficient fine-tuning (PEFT), and full fine-tuning. In former case, only a small subset of the model’s parameters are trainable, representing the performance in low computing resource available scenario. We adopt the most representative PEFT method LoRA [7] here. In latter case, all parameters will be fine-tuned which is a more conventional practice.
We have benchmarked the following LLMs on our proposed MMedBench:
  • GPT-4 [20], a groundbreaking multilingual large language model developed by OpenAI, stands as one of the most sophisticated LLMs to date. Due to the confidential of data and model details, it is uncertain for its detailed model scale. Though GPT-4 does not emphasize it is a multilingual LLM, its multilingual ability is still superior. Given that it is only accessible through API, we evaluate it in a zero-shot manner with proper instruction (Check Supplementary Material A for more details).
  • GPT-3.5 [1] is also developed by OpenAI. Similarly to GPT-4, it is unknown for detailed model sizes or training data composition and never claims whether it is a multilingual or monolingual LLM but it also performs well for multilingual input. As the predecessor to GPT-4, it continues to exhibit robust performance for everyday applications and remains extensively utilized. We assess GPT-3.5 using the same API-based methodology as applied to GPT-4.
  • Gemini-1.0 pro [24] is the latest general multimodal foundation model developed by Google. Though it is targeted at multimodal scenarios, as reported in the original paper, its language ability even surpasses Google’s former LLM, PaLM 2 [2]. Similar to GPT-series, its detailed scale and whether it is specifically targeted to multilingual or monolingual scenarios are not released. However, in our testing, it responds well for multilingual input.
  • BLOOM [22], an early open-source, multilingual LLM family, undergoes pre-training with a diverse range of language corpora. We select the 7B parameter variant for our studies, employing a fine-tuning evaluation approach.
  • MedAlpaca [6] is a specialized open-source monolingual medical LLM, further fine-tuned on a LLaMA using a dataset of over 160,000 English medical entries.
  • ChatDoctor [31] is a monolingual medical LLM based on LLaMA and further fine-tuned, leverages 100,000 real-world patient-doctor dialogues in English, marking it as a distinct medical LLM. We employ the 7B parameter model, applying a fine-tuning evaluation framework.
  • PMC-LLaMA [29] presents another open-source monolingual medical-specific LLM, pretrained exclusively on English medical literature, including papers and books. We utilize the 7B parameter version for evaluation.
  • LLaMA 2 [26], the latest iteration in the LLaMA series, is recognized as one of the most potent open-source monolingual LLMs with a focus on English. We engage the 7B parameter model in our finetuning evaluation process.
  • Mistral 7B [9], released in October 2023, is an innovative open-source monolingual LLM that claims superiority over LLaMA 2 13B across all assessed benchmarks. We adopt a fine-tuning evaluation methodology for this model.
  • InternLM and InternLM 2 [25], developed by Shanghai AI Lab, are among the leading open-source multilingual LLMs. InternLM was released in July 2023 and InternLM 2 was released in February 2024. For both models, we select the 7B parameter variant and implement a fine-tuning evaluation strategy.
4.3.5 Metrics and Human Rating
To evaluate the performance of LLMs, we employ two metrics: Accuracy and Rationale Similarity. Measuring Accuracy is straightforward, as the LLM can generate outputs following a specific template. However, assessing Rationale Similarity presents a more complex challenge, typical within the NLP domain. Initially, we applied three classical text similarity methods, BLEU [21] and ROUGE [16], and Bert-score [34].
BLEU quantifies the match between a model’s output and that of a reference, focusing on the precision of n-grams. The BLEU is calculated as follows:
$\text{BLEU}=\text{BP}\cdot\exp\left(\sum_{n=1}^{N}w_{n}\log P_{n}\right)$ (2)
where $P_{n}$ is the precision of n-grams, $w_{n}$ is the weight for each n-gram size, BP is the brevity penalty. $N$ typically equals 4 in most applications. To $\text{BLEU}-n$, the evaluation focuses only on n-grams of that specific length, by setting $w_{n}=1$ for a particular $n$ and setting all other weights to 0. In the standard BLEU calculation, a weighted average of BLEU-1 through BLEU-4 scores is used, with each component typically having equal weight($w_{1}=w_{2}=w_{3}=w_{4}=0.25$).
ROUGE is a metric that also focuses on n-grams but uniquely incorporates both Recall and Precision in its calculation. ROUGE is computed as follows:
$\text{ROUGE}=\frac{2\times P_{n}\times R_{n}}{P_{n}+R_{n}}$ (3)
where P and R represent Precision and Recall, respectively. Note that ROUGE-N emphasizes the Precision and Recall of n-grams, whereas ROUGE-L calculates the Precision and Recall based on the longest common subsequence (LCS).
BERT-score utilizes the contextual embedding from a pretrained BERT to capture high-level semantic features, calculating the similarity between reference and candidate texts through cosine similarity. The Recall is calculated as follows:
$R=\frac{\sum_{x_{i}\in x}\operatorname{idf}\left(x_{i}\right)\max_{\hat{x}_{j}\in\hat{x}}\mathbf{x}_{i}^{\top}\hat{\mathbf{x}}_{j}}{\sum_{x_{i}\in x}\operatorname{idf}\left(x_{i}\right)}$ (4)
where $\operatorname{idf}$ denotes inverse document frequency, enhancing the metric’s sensitivity to rare but significant words. Here, $x_{i}$ and $\hat{\mathbf{x}}_{j}$ represent the embeddings of the $i$th token in the candidate text and the $j$th token in the reference text, respectively. Precision is computed similarly, and the F1 score is subsequently derived. In this paper, we employ a pre-trained multilingual BERT model to extract features without conducting baseline rescaling.
Relative Rating scores aim to rank the output based on relative comparison. In detail, we select 6 representative models and sample 50 cases for each language. In human rating, for each case, question, options, right answer and rationale generated by each models are present to annotators, along with the reference rationale. The annotators are asked to rank the model-generated rationales based on the following three evaluation criteria:
  • Accuracy. The model’s ability to correctly select the answer.
  • Reasoning Ability. The model’s capacity to demonstrate logical reasoning in making its selection. The model should go beyond merely repeating the question or options, supporting its choice with reasonable reasoning.
  • Integration of Internal Knowledge. The model needs to effectively blend and utilize its internal knowledge base, providing relevant and persuasive factual evidence to support its answer.
Considering that GPT-4 has achieved near-human performance in many aspects, we use GPT-4 to rank the models in the same way with carefully set instructions following [35]. Similarly, for BLEU scores, we can also rank model by comparing the absolute metrics.
Then, for all ranking results, i.e., human rating, BLEU score rating and GPT-4 rating, the scores are quantitatively assigned with the ranking level reversely, for example, the top rank reviving a score of 6, and the bottom rank a score of 1, thereby quantifying each model’s output quality relatively.
4.4 Implementation Details
In this section, we delve into the specifics of auto-regressive training and fine-tuning. We conduct all our experiments using PyTorch framework and Transformers python package.
Auto-regressive Training. During further auto-regressive training on MMedC, our optimization objective aligns with that of the auto-regressive generation task. For data processing, we segment the text into chunks, each comprising 2048 tokens, with an overlapping margin of 512 tokens. Throughout the training, we maintain a maximum context length of 2048 tokens. Owing to the model’s extensive parameter count, which precludes fitting on a single GPU, we employ the Fully Sharded Data Parallel (FSDP) strategy to distribute the model across multiple GPU cards. Additionally, we utilize the BF16 data type and gradient checkpointing techniques to optimize memory usage. For InternLM, we establish a global batch size of 512 and a learning rate of 2e-5. In the case of BLOOM, we set the global batch size to 512 and a learning rate of 8e-6. We pretrain both models on eight A100 GPUs, adapting gradient accumulation steps to sustain such a large global batch size. We train the 7B model for 20k iterations, which takes about 20 Days.
Fine-tuning. During the fine-tuning process, our optimization objective remains consistent with the auto-regressive training phase. We set the maximum sequence length to 2048, padding each batch to match the longest sequence in that batch. For full-model fine-tuning, we utilized Fully Sharded Data Parallel (FSDP), BF16 data type, and gradient checkpointing technology. We set the global batch size to 128 and the learning rate to 1e-6. For LoRA, we use the default recommended rank 16 with the similar training setting as full fine-tuning.
5 Conclusion
In this paper, we describe an automatic pipeline for building up a new multilingual medical corpus together with a new benchmark. Specifically, we developed MMedC, a large-scale medical corpus with 25.5B tokens covering six main languages, and MMedBench, a comprehensive benchmark encompassing 6 primary languages. To monitor the development progress of medical multilingual LLMs, we have evaluated eleven existing LLMs on multi-choice question-answering and rationale generation abilities under various settings. Experimentally, we demonstrate the effectiveness of further training on MMedC, significantly filling the gap in adapting advanced general multilingual LLMs into complex and professional medical domain. As a result, we open-source the MMedLM 2, which to our knowledge, denotes the first, and strongest multilingual language model for medicine.
6 Data and Code Availability
Currently, we have released all of our data, code, and model weights.
Appendix A Used Prompts
In this section, we introduce the prompts used in our methodology and experiments.
A.0.1 Zero-shot prompt
For models such as GPT-3.5 and GPT-4, we employ specific instructions to steer the model’s response to questions. To prompt the model for a straightforward correct answer, we use the following directive: You’re a {language} doctor, make a choice based on the question and options. You need to answer the letter of the option instead of answering the entire option or anything else. Options may not be unique. When we aim for the model to provide not just the correct answer but also its corresponding rationale, we utilize a different instruction: You’re a {language} doctor, make a choice based on the question and options in {language}. You should solve this step-by-step. You must first give the reason in {language} for your choice ends with ’[End]’. Then you must give the answer’s letter directly again. The template is like ’Reason:… [End] Answer: A, B’
A.0.2 Fine-tuning Prompts
Similar to the prompting technique employed in GPT models, we utilize two distinct types of prompt instructions in our fine-tuning setting so that models can understand whether to output long rationale sentences or not, avoiding confusion when training on the mixture of rationale and multiple-choice data. In detail, for multiple-choice answering, we apply the following prompt: You’re a {language} doctor, kindly address the medical queries according to the patient’s account. Answer with the best option directly. In contrast, to obtain an answer accompanied by its corresponding rationale, we use a more detailed prompt: You’re a {language} doctor, kindly address the medical queries according to the patient’s account in {language}. Let’s solve this step-by-step. You should first give the reason in {language} for your choice. Then you should give the right answer index of the question. It is important to note that during fine-tuning, we employed both of the aforementioned instructions. This implies that for each sample, the model underwent optimization twice in each epoch, with different instructions and corresponding outputs. During evaluation, the first instruction was used to determine the model’s accuracy, owing to its significant convenience in metric calculation and the second instruction is utilized to generate rationales and to assess their similarity with the reference material.
A.0.3 Ranking Prompt
We used the following instructions for rating the rationales with GPT4. Please act as an impartial judge and evaluate the quality of the responses provided by six AI assistants to the user question displayed below. You should choose the assistant that follows the user’s instructions and answers the user’s questions better. Your evaluation should consider factors such as the helpfulness, relevance, accuracy, depth, creativity, and level of detail of their responses. Begin your evaluation by comparing the six responses. Avoid any position biases and ensure that the order in which the responses were presented does not influence your decision. Do not allow the length of the responses to influence your evaluation. Do not favor certain names of the assistants. Be as objective as possible. Your output is the ordering of these six models from high to low. Output your final verdict from high to low by strictly following this format: Model A, Model B, Model C, Model D, Model E, and Model F.
Appendix B MMedC Details
B.1 Settings for Filtering Content
In this section, we present the parameters employed for data filtering, which are meticulously configured based on the CulturaX dataset. Our parameterization adopts a relatively stringent criterion aimed at procuring high-quality data.
Table 4: Parameter Settings for Data Filtering. Given the extensive dataset within CulturaX, the configuration of filtering parameters is designed to prioritize quality over quantity.
[b]
Language Medical Keyword Count Keyword Density Remain Text(%) Tokens
English 5 4% 2.29% 6.56B
Spanish 4 4% 0.92% 3.98B
French 4 4% 0.50% 1.90B
Russian 4 2% 0.48% 1.29B
Chinese 5 5% 1.33% 3.34B
Japanese 5 5% 1.13% 1.93B
Table 4: Parameter Settings for Data Filtering. Given the extensive dataset within CulturaX, the configuration of filtering parameters is designed to prioritize quality over quantity.
B.2 Data Statistics
We give the details about the composition of MMedC, including the amount of data coming from different sources for each language.
Table 5: Corpus statistics. The Multilingual Medical Corpus (MMedC) boasts a substantial size of 105GB, comprising over 25 billion tokens. The dataset is categorized into several types by sources: FData represents filtering-based data, OCRData denotes Optical Character Recognition (OCR)-based data, CData refers to data obtained through web crawling, and OSData signifies data from open-source datasets. Furthermore, we distinguish between HQData (high-quality data) which has undergone human verification, and UQData (unsure-quality data), which has not. The TotAmt (Total Amount) quantifies the total token count for each language, with ’Billion tokens’ serving as the unit of measurement.
[b]
Language Family Filtering Content Textbooks Websites Small-scale Dataset HQData UQData TotAmt
English Indo-European 6.56 4.00 0.00 0.00 4.00 6.56 10.56
Spanish Indo-European 3.98 0.31 0.05 0.02 0.37 3.98 4.35
French Indo-European 1.90 0.02 0.00 0.17 0.20 1.90 2.10
Russian Indo-European 1.29 0.40 0.00 0.00 0.40 1.29 1.69
Chinese Sino-Tibetan 3.34 1.21 0.00 0.19 1.40 3.34 4.74
Japaneses Sino-Tibetan 1.93 0.00 0.10 0.01 0.12 1.93 2.05
Table 5: Corpus statistics. The Multilingual Medical Corpus (MMedC) boasts a substantial size of 105GB, comprising over 25 billion tokens. The dataset is categorized into several types by sources: FData represents filtering-based data, OCRData denotes Optical Character Recognition (OCR)-based data, CData refers to data obtained through web crawling, and OSData signifies data from open-source datasets. Furthermore, we distinguish between HQData (high-quality data) which has undergone human verification, and UQData (unsure-quality data), which has not. The TotAmt (Total Amount) quantifies the total token count for each language, with ’Billion tokens’ serving as the unit of measurement.
Appendix C Detailed Experiment Results
In this section, we present detailed results across languages in the ablation study, and rationale evaluation based on different metrics.
C.1 Detailed Accuracy on Ablation Study
Here, we give the detailed results of the ablation study, including both accuracy and rationale evaluation across six languages, as shown in Table 6 and 7.
[…]
Table 6: Accuracy across languages on ablation study.
Table 7: Rationale evaluation on MMedB with ROUGE-1/ BLEU-1. The value preceding the symbol ‘/’ represents BLEU-1, while the value following it represents ROUGE-1.
[b]
Method English Chinese Japanese French Russian Spanish Avg.
MMedLM
InternLM 46.53/41.86 48.24/48.64 44.89/49.83 41.80/37.95 27.87/21.20 43.42/38.59 42.12/39.68
[…]
MMedLM 47.37/41.98 48.68/49.28 48.95/52.34 45.39/41.41 33.24/24.67 46.68/41.35 45.05/41.84
MMedLM 2
InternLM2 49.48/44.12 51.38/51.58 50.64/53.46 46.73/42.00 32.93/24.05 47.94/41.96 46.52/42.86
MMedLM $\mathrm{2}_{en}$ 50.00/45.02 50.81/51.19 50.15/53.11 47.12/43.04 33.06/23.94 47.62/41.64 46.46/42.99
MMedLM $\mathrm{2}_{ocr}$ 49.82/44.75 51.47/52.19 52.66/55.21 47.84/43.68 36.98/28.17 48.93/43.42 47.95/44.57
MMedLM 2 50.02/44.77 51.39/51.78 54.79/57.10 49.04/45.30 37.49/28.18 50.14/44.59 48.81/45.29
Table 7: Rationale evaluation on MMedB with ROUGE-1/ BLEU-1. The value preceding the symbol ‘/’ represents BLEU-1, while the value following it represents ROUGE-1.
C.2 Different Metrics on Rationale Evaluation
We conducted an in-depth analysis of the scores assigned to the rationales generated by various models, utilizing a range of evaluation metrics. Table 8 and  9 illustrate the performance of each model as measured by the N-gram BLEU metrics under PEFT and full fine-tuning settings respectively. Meanwhile, Table 10 and 11 present the scores achieved by these models under the N-gram ROUGE metrics. Table 12 and 13 further show the BERT score of different models, which is a score calculated by averaging N-gram BERT scores as introduced in the main body.
[b] Method Metric English Chinese Japanese French Russian Spanish Avg. BLOOMZ BLEU-1 41.45 43.09 28.79 38.89 22.25 42.36 36.14 InternLM BLEU-1 41.29 43.47 22.80 30.35 18.24 36.32 32.08 LLaMA 2 BLEU-1 44.72 42.69 45.58 42.93 31.75 44.22 41.98 ChatDoctor BLEU-1 44.65 40.88 39.54 40.12 30.95 42.88 39.84 MedAlpaca BLEU-1 43.59 40.71 37.27 39.82 30.11 42.80 39.05 PMC-LLaMA BLEU-1 44.98 40.09 38.15 38.89 30.08 43.00 39.20 Mistral BLEU-1 48.13 45.61 43.82 44.73 33.62 47.37 43.88 InternLM 2 BLEU-1 46.87 47.64 42.22 41.81 26.78 44.51 41.64 MMedLM BLEU-1 41.63 44.30 38.61 37.54 19.99 40.79 37.14 MMedLM 2 BLEU-1 47.07 47.15 47.90 43.22 27.81 46.17 43.22 BLOOMZ BLEU-2 14.68 15.03 8.46 14.20 5.04 15.92 12.22 InternLM BLEU-2 14.79 15.28 7.29 9.92 4.59 11.59 10.58 LLaMA 2 BLEU-2 16.29 14.01 18.34 16.57 9.12 16.20 15.09 ChatDoctor BLEU-2 17.46 13.39 14.91 15.98 9.25 16.45 14.57 MedAlpaca BLEU-2 16.81 13.35 13.95 15.85 8.81 16.13 14.15 PMC-LLaMA BLEU-2 17.87 13.27 14.07 14.94 8.71 16.38 14.21 Mistral BLEU-2 20.33 16.26 17.11 18.56 10.56 19.32 17.02 InternLM 2 BLEU-2 18.91 18.53 16.29 15.80 7.22 16.53 15.55 MMedLM BLEU-2 15.33 16.06 14.69 13.76 5.96 15.12 13.49 MMedLM 2 BLEU-2 19.11 18.45 20.09 17.90 8.92 18.98 17.24 BLOOMZ BLEU-3 6.48 6.83 3.50 6.78 1.94 7.50 5.50 InternLM BLEU-3 6.64 6.97 3.48 4.42 1.80 4.93 4.71 LLaMA 2 BLEU-3 7.19 5.91 9.09 8.06 3.74 7.32 6.89 ChatDoctor BLEU-3 8.37 5.82 7.36 7.92 4.00 7.87 6.89 MedAlpaca BLEU-3 8.04 5.80 6.69 7.76 3.68 7.63 6.60 PMC-LLaMA BLEU-3 8.75 5.82 6.81 7.08 3.74 7.82 6.67 Mistral BLEU-3 10.35 7.40 8.57 9.61 4.56 9.68 8.36 InternLM 2 BLEU-3 9.12 8.98 8.10 7.69 2.84 7.83 7.42 MMedLM BLEU-3 6.92 7.48 7.59 6.44 2.33 7.11 6.31 MMedLM 2 BLEU-3 9.21 8.99 10.59 9.21 3.87 9.75 8.60 BLOOMZ BLEU-4 3.25 3.38 1.77 3.26 0.87 4.02 2.76 InternLM BLEU-4 3.53 3.41 1.87 2.14 0.91 2.49 2.39 LLaMA 2 BLEU-4 3.72 2.76 4.94 4.06 1.77 3.85 3.52 ChatDoctor BLEU-4 4.52 2.62 3.99 4.10 1.95 4.37 3.59 MedAlpaca BLEU-4 4.50 2.72 3.60 3.96 1.74 4.18 3.45 PMC-LLaMA BLEU-4 4.90 2.73 3.71 3.46 1.76 4.30 3.48 Mistral BLEU-4 5.92 3.75 4.72 5.18 2.16 5.63 4.56 InternLM 2 BLEU-4 5.07 4.74 4.41 4.07 1.25 4.21 3.96 MMedLM BLEU-4 3.70 3.77 4.27 3.16 1.08 3.87 3.31 MMedLM 2 BLEU-4 5.21 4.71 5.96 5.03 1.93 5.61 4.74 BLOOMZ BLEU 9.95 10.29 5.61 9.91 3.19 11.32 8.38 InternLM BLEU 10.34 10.43 5.33 6.76 2.95 7.79 7.27 LLaMA 2 BLEU 11.14 9.18 13.46 11.69 5.89 11.18 10.42 ChatDoctor BLEU 12.50 8.84 10.94 11.30 6.14 11.87 10.26 MedAlpaca BLEU 12.16 8.93 10.15 11.17 5.76 11.56 9.95 PMC-LLaMA BLEU 13.07 8.91 10.31 10.28 5.80 11.71 10.01 Mistral BLEU 15.10 11.34 12.69 13.59 6.94 14.27 12.32 InternLM 2 BLEU 13.55 13.27 12.05 11.43 4.36 11.76 11.07 MMedLM BLEU 10.66 11.21 11.17 9.51 3.58 10.68 9.47 MMedLM 2 BLEU 13.84 13.16 15.21 13.06 5.73 14.21 12.54
Table 8: Rationale evaluation of models undergo PEFT using n-grams BLEU metrics.
[b] Method Metric English Chinese Japanese French Russian Spanish Avg. BLOOMZ BLEU-1 45.94 48.37 44.71 44.47 29.95 45.91 43.22 InternLM BLEU-1 46.53 48.24 44.89 41.80 27.87 43.42 42.12 LLaMA 2 BLEU-1 46.87 46.62 48.53 44.43 33.05 45.96 44.24 ChatDoctor BLEU-1 47.22 44.66 38.87 44.64 32.19 45.68 42.21 MedAlpaca BLEU-1 47.33 45.72 45.35 43.78 32.80 45.99 43.49 PMC-LLaMA BLEU-1 47.33 45.87 44.52 43.80 31.14 46.30 43.16 Mistral BLEU-1 47.16 48.34 48.80 45.83 34.52 47.55 45.37 InternLM 2 BLEU-1 49.48 51.38 50.64 46.73 32.93 47.94 46.52 MMedLM BLEU-1 47.37 48.68 48.95 45.39 33.24 46.68 45.05 MMedLM 2 BLEU-1 50.02 51.39 54.79 49.04 37.49 50.14 48.81 BLOOMZ BLEU-2 17.93 18.70 17.89 18.29 8.58 17.69 16.51 InternLM BLEU-2 18.76 18.79 18.55 15.89 7.44 15.89 15.89 LLaMA 2 BLEU-2 18.53 16.98 20.42 17.72 9.40 17.31 16.73 ChatDoctor BLEU-2 19.50 16.29 15.36 18.25 9.76 17.87 16.17 MedAlpaca BLEU-2 19.30 16.41 18.48 17.82 9.90 17.91 16.64 PMC-LLaMA BLEU-2 20.05 16.28 17.67 17.99 9.10 17.68 16.46 Mistral BLEU-2 18.57 17.97 20.47 19.02 10.48 18.33 17.47 InternLM 2 BLEU-2 21.20 20.76 22.83 20.14 9.72 19.12 18.96 MMedLM BLEU-2 19.39 18.84 21.63 19.46 10.13 18.99 18.07 MMedLM 2 BLEU-2 21.52 21.46 26.75 23.18 13.42 21.54 21.31 BLOOMZ BLEU-3 8.39 8.97 8.90 9.15 3.66 8.42 7.92 InternLM BLEU-3 9.20 9.05 9.56 7.56 3.09 7.24 7.62 LLaMA 2 BLEU-3 8.89 7.64 10.41 8.78 3.63 7.98 7.89 ChatDoctor BLEU-3 9.62 7.34 7.74 9.17 4.10 8.63 7.77 MedAlpaca BLEU-3 9.62 7.66 9.39 8.98 4.16 8.65 8.08 PMC-LLaMA BLEU-3 10.14 7.27 8.91 8.91 3.88 8.46 7.93 Mistral BLEU-3 8.67 8.42 10.33 9.65 4.41 8.76 8.37 InternLM 2 BLEU-3 10.71 10.54 12.41 10.55 4.12 9.31 9.61 MMedLM BLEU-3 9.55 9.20 11.84 10.24 4.20 9.38 9.07 MMedLM 2 BLEU-3 11.00 11.02 15.05 12.99 6.37 11.22 11.27 BLOOMZ BLEU-4 4.55 4.62 4.87 4.73 1.76 4.58 4.18 InternLM BLEU-4 5.23 4.82 5.51 3.88 1.55 3.90 4.15 LLaMA 2 BLEU-4 4.85 3.67 5.94 4.66 1.66 4.28 4.17 ChatDoctor BLEU-4 5.47 3.65 4.42 4.86 1.96 4.71 4.18 MedAlpaca BLEU-4 5.43 3.88 5.33 4.83 1.95 4.89 4.38 PMC-LLaMA BLEU-4 5.79 3.58 5.04 4.52 1.93 4.66 4.25 Mistral BLEU-4 4.75 4.30 5.74 5.16 2.18 4.88 4.50 InternLM 2 BLEU-4 6.13 5.84 7.41 5.83 2.03 5.17 5.40 MMedLM BLEU-4 5.42 5.02 7.09 5.71 1.88 5.39 5.08 MMedLM 2 BLEU-4 6.41 6.20 9.15 7.70 3.30 6.63 6.56 BLOOMZ BLEU 12.80 13.14 13.21 13.06 5.70 12.53 11.74 InternLM BLEU 13.73 13.46 13.95 11.19 4.97 11.17 11.41 LLaMA 2 BLEU 13.32 11.44 15.28 12.72 5.83 12.10 11.78 ChatDoctor BLEU 14.10 11.09 11.49 13.09 6.31 12.81 11.48 MedAlpaca BLEU 14.05 11.49 13.74 12.99 6.36 12.97 11.94 PMC-LLaMA BLEU 14.74 11.01 13.22 12.76 6.06 12.77 11.76 Mistral BLEU 13.11 12.65 15.16 13.77 6.90 13.25 12.47 InternLM 2 BLEU 15.44 15.28 17.59 14.90 6.48 13.80 13.91 MMedLM BLEU 14.01 13.69 16.72 14.38 6.36 13.80 13.16 MMedLM 2 BLEU 15.92 15.80 20.75 17.72 9.35 16.11 15.94
Table 9: Rationale evaluation of models undergo full fine-tuning using n-grams BLEU metrics.
Table 10: Rationale evaluation of models undergo PEFT using n-grams ROUGE metrics.
[b]
Method Metric English Chinese Japanese French Russian Spanish Avg.
BLOOMZ ROUGE-1 36.81 45.17 38.09 37.49 15.28 39.22 35.34
InternLM ROUGE-1 38.05 44.71 37.57 32.14 16.79 34.56 33.97
LLaMA 2 ROUGE-1 39.34 43.71 49.53 39.29 22.66 39.64 39.03
ChatDoctor ROUGE-1 40.26 42.80 45.00 39.06 22.84 40.23 38.37
MedAlpaca ROUGE-1 39.52 42.50 44.69 39.57 22.83 39.64 38.12
PMC-LLaMA ROUGE-1 40.90 42.95 43.67 38.64 22.45 39.80 38.07
Mistral ROUGE-1 42.80 46.31 48.19 41.07 24.75 42.83 40.99
InternLM 2 ROUGE-1 41.66 49.28 46.91 38.46 21.71 40.13 39.69
MMedLM ROUGE-1 38.83 46.38 46.90 37.78 21.28 38.77 38.32
MMedLM 2 ROUGE-1 41.51 48.36 52.24 41.36 25.70 42.64 41.97
BLOOMZ ROUGE-2 12.73 16.77 13.43 14.01 4.03 15.66 12.77
InternLM ROUGE-2 13.08 16.71 13.92 10.80 4.43 12.23 11.86
LLaMA 2 ROUGE-2 13.63 15.18 21.37 15.06 6.41 15.13 14.47
ChatDoctor ROUGE-2 15.25 15.10 19.18 15.64 6.82 16.13 14.69
MedAlpaca ROUGE-2 14.77 15.09 18.86 15.68 6.65 15.83 14.48
PMC-LLaMA ROUGE-2 15.59 15.37 18.11 14.81 6.66 15.90 14.41
Mistral ROUGE-2 17.28 17.60 20.49 17.00 7.72 18.10 16.36
InternLM 2 ROUGE-2 15.95 20.04 20.26 14.55 6.00 15.62 15.40
MMedLM ROUGE-2 13.78 17.68 19.75 14.44 6.22 15.27 14.52
MMedLM 2 ROUGE-2 16.02 19.94 23.27 17.01 8.29 18.22 17.12
BLOOMZ ROUGE-L 33.28 30.71 26.91 34.01 14.56 35.22 29.11
InternLM ROUGE-L 34.26 31.02 26.31 29.37 15.76 30.99 27.95
LLaMA 2 ROUGE-L 34.91 28.40 31.82 35.38 21.02 34.99 31.08
ChatDoctor ROUGE-L 36.57 28.64 30.86 35.63 21.15 35.98 31.47
MedAlpaca ROUGE-L 35.88 28.87 30.35 36.08 21.47 35.41 31.34
PMC-LLaMA ROUGE-L 36.84 28.90 29.43 34.96 20.92 35.35 31.07
Mistral ROUGE-L 38.57 31.23 31.49 37.23 23.18 37.90 33.27
InternLM 2 ROUGE-L 37.45 34.57 31.70 34.95 20.23 35.78 32.45
MMedLM ROUGE-L 35.08 32.18 32.13 34.19 19.70 34.41 31.28
MMedLM 2 ROUGE-L 37.61 33.52 33.92 37.32 23.95 37.95 34.04
Table 10: Rationale evaluation of models undergo PEFT using n-grams ROUGE metrics.
[b] Method Metric English Chinese Japanese French Russian Spanish Avg. BLOOMZ ROUGE-1 40.51 48.26 48.61 41.05 21.50 40.77 40.12 InternLM ROUGE-1 41.86 48.64 49.83 37.95 21.20 38.59 39.68 LLaMA 2 ROUGE-1 41.39 46.57 51.21 40.38 23.24 40.37 40.53 ChatDoctor ROUGE-1 41.97 45.81 47.95 40.25 23.37 40.71 40.01 MedAlpaca ROUGE-1 42.31 46.49 49.12 40.41 23.15 40.57 40.34 PMC-LLaMA ROUGE-1 42.87 46.18 48.44 40.23 22.28 40.68 40.12 Mistral ROUGE-1 41.82 47.91 50.60 40.88 24.68 41.41 41.22 InternLM 2 ROUGE-1 44.12 51.58 53.46 42.00 24.05 41.96 42.86 MMedLM ROUGE-1 41.98 49.28 52.34 41.41 24.67 41.35 41.84 MMedLM 2 ROUGE-1 44.77 51.78 57.10 45.30 28.18 44.59 45.29 BLOOMZ ROUGE-2 15.25 19.78 21.22 16.63 6.13 16.44 15.91 InternLM ROUGE-2 16.19 19.86 22.13 14.67 5.64 14.92 15.57 LLaMA 2 ROUGE-2 15.75 18.07 22.81 15.76 6.33 15.81 15.76 ChatDoctor ROUGE-2 16.70 17.67 20.65 16.16 7.12 16.64 15.82 MedAlpaca ROUGE-2 16.41 17.63 21.78 16.21 7.11 16.49 15.94 PMC-LLaMA ROUGE-2 17.40 17.56 21.00 16.20 6.77 16.47 15.90 Mistral ROUGE-2 15.64 18.71 22.64 16.64 7.31 16.71 16.27 InternLM 2 ROUGE-2 18.04 21.75 25.32 17.76 7.02 17.36 17.88 MMedLM ROUGE-2 16.31 20.00 24.50 17.82 7.57 17.56 17.29 MMedLM 2 ROUGE-2 18.47 22.56 28.88 20.80 10.02 19.68 20.07 BLOOMZ ROUGE-L 36.53 32.84 32.03 37.17 19.90 35.94 32.40 InternLM ROUGE-L 37.89 33.59 32.49 34.37 20.13 34.26 32.12 LLaMA 2 ROUGE-L 37.05 31.09 32.45 36.44 21.73 35.72 32.41 ChatDoctor ROUGE-L 37.91 31.04 31.62 36.29 22.02 36.38 32.54 MedAlpaca ROUGE-L 38.00 31.43 32.08 36.57 21.77 36.13 32.66 PMC-LLaMA ROUGE-L 38.77 30.85 31.42 36.57 21.13 36.19 32.49 Mistral ROUGE-L 37.22 32.06 31.78 36.88 23.00 36.35 32.88 InternLM 2 ROUGE-L 39.61 36.27 35.06 38.17 22.74 37.24 34.85 MMedLM ROUGE-L 37.84 34.35 34.51 37.81 23.19 37.08 34.13 MMedLM 2 ROUGE-L 40.19 36.97 38.31 41.17 26.62 39.71 37.16
Table 11: Rationale evaluation of models undergo full fine-tuning using n-grams ROUGE metrics.
[b] Method English Chinese Japanese French Russian Spanish Avg. BLOOMZ 74.72 76.46 67.87 76.36 67.24 75.58 73.04 InternLM 75.13 76.95 67.22 72.46 67.87 72.94 72.09 LLaMA 2 76.50 76.38 74.81 77.77 74.30 76.54 76.05 ChatDoctor 76.53 75.23 72.31 77.34 74.25 76.66 75.39 MedAlpaca 76.30 75.43 71.33 77.36 74.14 76.43 75.16 PMC-LLaMA 76.84 75.17 71.28 77.15 73.77 76.35 75.09 Mistral 77.99 77.46 74.11 78.67 75.07 77.77 76.85 InternLM 2 77.36 78.75 73.05 77.58 73.68 76.66 76.18 MMedLM 75.68 77.16 72.15 76.41 73.87 75.42 75.12 MMedLM 2 77.59 78.42 75.79 78.47 76.07 77.65 77.33
Table 12: Rationale evaluation of models undergo PEFT using BERT Score metrics.
[b] Method English Chinese Japanese French Russian Spanish Avg. BLOOMZ 76.66 78.30 74.26 78.18 72.22 76.92 76.09 InternLM 77.11 78.43 74.73 77.04 71.91 76.04 75.88 LLaMA 2 77.49 77.78 75.99 78.45 74.60 77.21 76.92 ChatDoctor 77.47 77.18 72.65 78.16 74.69 77.19 76.22 MedAlpaca 77.53 77.31 74.39 78.23 74.85 77.24 76.59 PMC-LLaMA 77.98 77.22 74.15 78.15 73.94 76.96 76.40 Mistral 77.79 78.49 76.00 79.00 75.77 77.80 77.47 InternLM 2 78.44 79.37 76.70 79.38 74.95 78.01 77.81 MMedLM 77.58 78.66 76.02 78.71 75.42 77.36 77.29 MMedLM 2 78.74 79.56 78.40 80.36 77.41 78.73 78.87
Table 13: Rationale evaluation of models undergo full fine-tuning using BERT Score metrics.
C.3 Detailed Results on Rating Scores
In Table 14 and 15, we further show the GPT-4 and human rating scores for six languages respectively.
Table 14: Detailed scores of GPT-4 rating. The scores are quantitatively assigned with the ranking level reversely.
[b]
Method English Chinese Japanese French Russian Spanish Avg.
MMedLM 4.12 4.74 4.44 4.44 4.22 4.36 4.39
InternLM 3.58 4.04 3.26 2.74 2.64 2.8 3.18
BLOOMZ 2.68 3.18 3.02 3.2 3.28 3.22 3.1
LLaMA 2 3.24 3.2 3.46 3.2 4.18 3.72 3.5
PMC-LLaMA 4.3 3.22 3.06 3.22 3.26 3.2 3.38
GPT-3.5 3.08 2.62 3.76 4.2 3.42 3.7 3.46
Table 14: Detailed scores of GPT-4 rating. The scores are quantitatively assigned with the ranking level reversely.
[b] Method English Chinese Japanese French Russian Spanish Avg. MMedLM 4.3 4.78 4.94 4.14 4.44 4.66 4.54 InternLM 3.68 3.94 3.78 3.44 2.92 2.62 3.4 BLOOMZ 3.2 3.6 3.16 3.3 3.34 3.26 3.31 LLaMA 2 3.14 3.48 3.52 3.5 3.68 3.96 3.55 PMC-LLaMA 3.44 2.72 2.82 3.2 3.38 3.16 3.12 GPT-3.5 3.24 2.48 2.78 3.42 3.24 3.34 3.08
Table 15: Detailed scores of human rating. The scores are quantitatively assigned with the ranking level reversely.
Appendix D Case Study
In this section, we present an analysis of individual cases across various languages. We first analyze the output of models in a zero-shot setting, and then we analyze the output of models in a full fine-tuning setting.
D.1 Zero-shot Setting
In a zero-shot setting, we observe that it is challenging for a 7B-scale language model to consistently follow instructions, making it difficult to calculate accuracy automatically. Though difference cases, we notice that even when given the same instruction of "Answer with the best option directly.", the model will still express it in many ways. ###The most appropriate next step in management for this patient is to perform a bone marrow biopsy. (…) The best treatment for this patient is option C. (…) Given the patient’s symptoms and medical history, the most likely diagnosis is HIV infection. (…) The most appropriate next step in management for this patient is D. Ruxolitinib. (…)
What’s more, we find that in a zero-shot setting, the model is very sensitive to the instructions. For instance, only changing the instruction to "Answer the letter of the option directly" may lead the model to generate a different answer. Considering the difficulty in finding a fair way to automatically evaluate models’ performance in zero-shot setting, we turn to fine-tuning settings. The most appropriate next step in management for this patient is D. Ruxolitinib. This is because the patient’s symptoms (…) B### Explanation:The patient’s symptoms and laboratory findings suggest a myeloproliferative neoplasm (…)
D.2 Full Fine-tuning Setting
We further specifically compared the responses of MMedLM and InternLM, noting that MMedLM represents an advanced version of InternLM, augmented with domain-specific knowledge. As illustrated in Figures 6, 7, 8, 9, 10, and 11, MMedLM consistently demonstrates superior performance in selecting the correct option, unlike InternLM. Furthermore, MMedLM frequently justifies its choices with accurate reasons. For instance, in Figure 6, MMedLM accurately diagnoses ‘eosinophilic infiltration in the renal biopsy,’ subsequently applying its domain knowledge to identify these findings as indicative of toxic tubulointerstitial nephritis, leading to a precise diagnosis.
Nevertheless, we also notice an exception in Figure 10, where the language model (LLM) correctly identifies the answer, but with an erroneous rationale. Although the model attributes apoptosis to viral infections by stating that ’viruses attack cells, causing apoptosis,’ it overlooks the fact that apoptosis is, in reality, an active defense mechanism employed by the body to curb viral replication. This observation suggests that the model’s selection of the correct answer could be attributed to chance rather than a genuine understanding or possession of relevant knowledge. It further underscores the importance of evaluating not only the accuracy of the model’s answers but also its reasoning process.
Figure 6: Case 1. An English case in MMedBench
Figure 7: Case 2. A Chinese case in MMedBench
Figure 8: Case 3. A Japanese case in MMedBench
Figure 9: Case 4. A French case in MMedBench
Figure 10: Case 5. A Russian case in MMedBench
Figure 11: Case 6. A Spanish case in MMedBench